### v1

@@ IV-B ${\tt oSCC}$ Planner
[…]
@@
Figure 6: An illustrative planning example by the ${\tt oSCC}$ algorithm. (a)-(c) The Reeb graphs as additional cracks are detected during the loop cycles of the ${\tt oSCC}$ algorithm. All cracks are marked by thick, dotted curves. The robot starts from the top-left corner. Each Reeb graph is updated when the robot is at point $A$ (highlighted in a circle). Those Reeb graphs are constructed after removing already-covered areas. The red dots ($C_{i}$) denote the nodes of the Reeb graph. The solid curves ($E_{i}$) are the edges. The doubled edges are indicated by the thin, dashed black curves ($E_{i}^{\prime}$). The bold boundaries are the cell boundaries. The orange-dash lines are the robot paths. The blue-dotted dashed lines are the connections between two nodes. (d) The final path using the ${\tt oSCC}$ algorithm. The final MCD of the free space is plotted in the thick orange curves. The final trajectory is shown by the black dashed lines. The robot starts from the top-left corner and ends at the bottom left, marked by a triangle and an arrow, respectively.
In Algorithm 4, after traversing one target edge, the covered areas and update $\mathbb{G}_{\text{w}}$ of the remaining space $\mathcal{W}\backslash\mathcal{M}_{\text{cov}}$ are removed to avoid passing the same target twice to reach another uncovered cell. Taking the example shown in Fig. 6, the robot is currently on node $A$ and $\mathbb{G}_{\text{w}}$ is updated after removing the already covered areas (yellow shaded areas). The odd nodes in the resulting $\mathbb{G}_{\text{w}}$ include $C_{1}$, $C_{2}$, $C_{3}$, and $C_{4}$. Therefore, $AC_{2}$ and $C_{1}C_{3}$ are connected to form the least-cost Euler tour, and $C_{4}$ represents the path’s ending node. When the resulting Euler tour needs to double edges in $\mathbb{G}_{\text{w}}$, the corresponding cell is split into two components. The first part is covered by a wall-following motion, where a wall is defined as the boundaries of $\mathcal{W}\backslash\mathcal{M}_{\text{cov}}$. The other part is covered by the zigzag motion of the leftover space in the cell. As shown in Fig. 6, edges $E_{1}$ and $E_{2}$ are doubled, and in Fig. 6, edges $E_{2}$ and $E_{3}$ are doubled. Note that the splitting of cells does not increase the cost of covering the whole cell. To minimize the cost, we select the coverage motion direction according to the next connected edge in the path.
@@
The ${\tt oSCC}$ algorithm ensures that each individual cell in the free space is covered exactly once, avoiding redundancy in terms of individual cell coverage when the connections of each covered cell are not considered. Because the dimensions of the cells are unknown, the robot does not always have enough information to minimize the zigzag motion connecting adjacent cells according to the Euler tour sequence. The optimality gap in our approach is related to the connections between adjacent cells based on the Euler tour sequence. For a worst-case analysis of the suboptimality bound, we assume that the connections between each pair of adjacent cells are small enough to be negligible in the optimal solution, and the ${\tt oSCC}$ result provides the worst connection between each pair of adjacent cells, which is the number of free cells in $\mathbb{G}_{\text{w}}$, denoted as $N_{\text{cell}}$, minus one (i.e., the number of connections), multiplied by $\sqrt{l^{2}+(2S)^{2}}$ (i.e., the longest robot travel distance for one connection), where $l$ is the slice length (the side length of the free space) and $S$ is the sensing radius. Therefore, in the worst case, the difference from the optimal solution is bounded by $(N_{\text{cell}}-1)\sqrt{l^{2}+(2S)^{2}}$. The simplified $\mathbb{G}_{\text{w}}$ reduces the number of edges. The algorithm is solved in polynomial time because of the structure of $\mathbb{G}_{\text{w}}$ and is therefore used online efficiently.
We further illustrate the ${\tt oSCC}$ and ${\tt SCC}$ planners through an example. Fig. 6 shows the planning result under ${\tt oSCC}$ in $\mathcal{W}$ with sizes $l=5.79$ m and $w=6.10$ m and robot configuration, $S=0.69$ m and $a=8.9$ cm. The crack information (thick dotted lines) is initially unknown. Under ${\tt oSCC}$ planner, the robot first follows the initial path planned by Algorithm 4 until detecting a crack node. Then it follows the online updated $\mathbb{G}_{\text{c}}$ and uses the ${\tt GCC}$ algorithm to generate local paths for footprint coverage of the scanned cracks. The covered regions are then removed from $\mathcal{W}$, and $\mathbb{G}_{\text{c}}$ and $\mathbb{G}_{\text{w}}$ are updated accordingly. Figs. 6-6 show the updated graphs, and Fig. 6 shows the final path. For comparison, Fig. 7 shows the planning results under the ${\tt SCC}$ algorithm with known crack information. The constructed MCD of the free space is shown in Fig. 7, along with the corresponding $\mathbb{G}_{\text{w}}$ (solid lines), $\mathbb{G}_{\text{c}}$ (dash lines), and $\mathcal{M}_{\text{c}}$ (yellow shaded areas). The least-cost Euler tour is constructed by the blue dotted dash lines. Fig. 7 shows the final route of the robot. Readers can refer to the companion video clip for details. Experiments and comparisons will be further discussed in Section VII.
@@
Figure 7: Planning result by the ${\tt SCC}$ algorithm with the cracks as in Fig. 6. (a) The MCD and the corresponding $\mathbb{G}_{\text{w}}$ using ${\tt SCC}$. The orange dotted lines are the constructed $\mathbb{G}_{\text{c}}$. The blue-dotted dashed lines show the node connections. (b) The final route of the robot. The legends are the same as those in Fig. 6.
@@ V-B Filling Nozzle Motion Planning
The nozzle motion needs to be coordinated with the robot motion to efficiently fill cracks within $\mathcal{F}$; see Fig. 8. Note that multiple cracks can be located within $\mathcal{F}$ at a time, and we need to determine how to move the nozzle to fill these cracks while the robot is in motion. Compared to robot moving velocity, the nozzle motion in general is much faster, and therefore, for simplicity, we neglect the time duration for the nozzle to move from one crack to another (no filling action) and only count movement time along cracks (with filling action).
Fig. 9 illustrates the geometric relationship between robot trajectory $\mathcal{P}_{R}$ and multiple cracks within $\mathcal{F}$. We use the arc length of path $\mathcal{P}_{R}$, denoted as $s$, as the parameter to characterize any arbitrary point ${p}$ on cracks. For point $p$ on a crack within $\mathcal{F}$, we define a projection map $\pi(p):\;{p}\mapsto p_{c}(s)$, $p_{c}(s)\in\mathcal{P}_{R}$, as the minimum distance to $p$. Considering that $n_{c}$ cracks are located within $\mathcal{F}$, $n_{c}\in\mathbb{N}$, we denote mapping $\pi_{i}(p)$ for the $i$th crack by the above definition, $i=1,\cdots,n_{c}$. We assume that $\pi_{i}(p)$ is bijective so that its inverse $\pi^{-1}_{i}(p)$ exists. Using $\pi_{i}(p)$, all points on $n_{c}$ cracks inside $\mathcal{F}$ are mapped onto $\mathcal{P}_{R}$.
@@
Value $s$ increases along the robot $\mathbf{R}$’s moving direction, and points on $n_{c}$ cracks within $\mathcal{F}$ with the same $s$ value have the same priority to fill. To determine the filling sequences, we consider the nozzle $\mathbf{N}_{z}$ to move and switch the filling action among these $n_{c}$ cracks after staying along one crack for a threshed distance $L$. Fig. 9 illustrates the nozzle travel sequence among all cracks within $\mathcal{F}$. Algorithm 5 describes the nozzle motion sequence planning. Input $\{{\boldsymbol{p}}_{i}\}_{1}^{n_{c}}$ is the point sequence sets of all cracks; that is, $\{{\boldsymbol{p}}_{i}\}$ contains the point sequence of the $i$th crack, $i=1,\cdots,n_{c}$. ${\tt startpoint}$ (line 1) is to find the first point of each crack. The function ${\tt findmin}$ is to find out the minimum value and the corresponding crack index. The notation $h(i)/h(I_{1})$ indicates removing $h(I_{1})$ from all $h(i)$ (line 3). The function ${\tt addpoint}$ (line 4) is to add an interval point of the crack $I_{1}$ to ${}^{s}{\boldsymbol{p}}_{n}$, and the interval is from $p_{1}$ to $p_{2}+L$. Finally, we obtain the output $\{{\boldsymbol{p}}_{n}\}$ that represents the point sequence of the nozzle in $\mathcal{B}$.
@@
21 ${\boldsymbol{p}}_{c}^{i}\leftarrow\pi_{i}({\boldsymbol{p}}_{i})$, $h(i)\leftarrow{\tt startpoint}({\boldsymbol{p}}_{c}^{i})$, $i=1,\cdots,n_{c}$ $(p_{1},I_{1})\leftarrow{\tt findmin}(h(i))$, $m\leftarrow 0$ while $m<n_{c}$ do
      76543 $(p_{2},I_{2})\leftarrow{\tt findmin}[h(i)/h(I_{1})]$ ${}^{s}{\boldsymbol{p}}_{m}\leftarrow{\tt addpoint}({\boldsymbol{p}}_{c}(I_{1}),p_{1},p_{2}+L)$ if ${\tt isfinish}(I_{1})$ then $m\leftarrow m+1$, $h(I_{1})\leftarrow inf$ else  $h(I_{1})\leftarrow p_{2}+L$ $(p_{1},I_{1})\leftarrow(p_{2},I_{2})$
@@
As shown in Fig. 9, by Algorithm 5, the first points for each crack of the first three cracks are $h(1)=s_{1a}$, $h(2)=s_{2a}$, and $h(3)=s_{3a}$. In the algorithm, $p_{1}=s_{1a}$ and $I_{1}=1$ since $s_{1a}$ is the minimum point. Then, we find the minimum from $h(2)$ and $h(3)$, so that $p_{2}=s_{2a}$, $I_{2}=2$ (line 3). We also obtain $s_{1b}=s_{2a}+L$ (line 4), and therefore, all points on crack 1 from $s_{1a}$ to $s_{1b}$ are added to ${}^{s}{\boldsymbol{p}}_{1}$. The next iteration then adds the points on crack 2 and keeps going until all $n_{c}$ cracks are covered. The value of $L$ is chosen to provide a trade-off between the nozzle switching frequency among cracks and the dedicating time duration to a single crack. When $L$ is small, the switching between different cracks becomes frequent, and when $L$ is large, it might miss filling some cracks within $\mathcal{F}$. We consider using $L\leq 2a$ for the switching between cracks within $\mathcal{F}$ to have a good performance.
@@ V-C Robot and Nozzle Motion Control
The robot’s orientation is independently controlled with linear velocity due to the use of omni-directional wheels. We first show the choice of orientation control. The robot driving energy expense can be represented as $J_{\theta}=\frac{1}{2}\sum_{i=1}^{4}\omega_{i}^{2}$. Taking the derivative of $J_{\theta}$ over $\theta$, we obtain
@@
Using (1), it is straightforward to show that $\frac{\partial J_{\theta}}{\partial\theta}=0$ and the energy expense is independent of the robot orientation. Similarly, we obtain
@@
The above result implies that changing the orientation increases the energy cost. Therefore, from (3) and (4), we set $\dot{\theta}=0$ to minimize the energy expense $J_{\theta}$, and in implementation, we further keep $\theta=0$ as the desired body orientation during robot movement for simplicity.
For the robot and nozzle motion control, the objective is to reach all cracks within $\mathcal{F}$ and complete the crack-filling action while the robot $\mathbf{R}$ is in motion. The desired nozzle path is given by Algorithm 5 in $\mathcal{B}$ as $\{{\boldsymbol{p}}_{n}\}$, while the desired robot path $\mathcal{P}_{R}$ is denoted as ${\boldsymbol{p}}_{r}$ in $\mathcal{N}$. $\mathcal{P}_{R}$ is computed from the coverage path planning algorithms, specifically Algorithms 3, 4, and 6. Defining ${\boldsymbol{\xi}}=[{\boldsymbol{q}}_{r}^{\mathsf{T}}\;{\boldsymbol{q}}_{n}^{\mathsf{T}}]^{\mathsf{T}}$ and ${\boldsymbol{u}}=[{\boldsymbol{\omega}}_{r}^{\mathsf{T}}\;{\boldsymbol{\omega}}_{n}^{\mathsf{T}}]^{\mathsf{T}}$, from (1), a discrete-time state-space model is used to represent the robot and nozzle motions at the $k$th step
@@
The robot velocity is slow compared with the nozzle motion. We denote $\mathcal{X}_{r}\subset\mathbb{R}^{3}$ and $\mathcal{X}_{n}\subset\mathbb{R}^{2}$ as the allowable robot and filling nozzle velocity sets in $\mathcal{N}$, respectively. We then have the velocity constraints as $\|{\boldsymbol{v}}_{r}(k)\|_{2}\leq\|{\boldsymbol{v}}_{n}(k)\|_{2}$, where ${\boldsymbol{v}}_{r}(k)=\dot{{\boldsymbol{q}}}_{r}(k)\in\mathcal{X}_{r}$ and ${\boldsymbol{v}}_{n}(k)=\dot{{\boldsymbol{q}}}_{n}(k)\in\mathcal{X}_{n},$ are the robot velocity and the nozzle relative velocity in $\mathcal{N}$, respectively. Using (1), ${\boldsymbol{v}}_{r}$ and ${\boldsymbol{v}}_{n}$ are calculated as
@@
$v_{n1}$ $=$ $\frac{R_{g}}{2}(\omega_{5}+\omega_{6})\mathop{\rm c}\nolimits_{\theta}-{{}^{b}x_{n}}\dot{\theta}\mathop{\rm s}\nolimits_{\theta}+\frac{R_{g}}{2}(\omega_{5}-\omega_{6})\mathop{\rm s}\nolimits_{\theta}-$
${{}^{b}y_{n}}\dot{\theta}\mathop{\rm c}\nolimits_{\theta}+\dot{x}_{r},$
$v_{n2}$ $=$ $\frac{R_{g}}{2}(\omega_{5}+\omega_{6})\mathop{\rm s}\nolimits_{\theta}+{{}^{b}x_{n}}\dot{\theta}\mathop{\rm c}\nolimits_{\theta}-\frac{R_{g}}{2}(\omega_{5}-\omega_{6})\mathop{\rm c}\nolimits_{\theta}-$
${{}^{b}y_{n}}\dot{\theta}\mathop{\rm s}\nolimits_{\theta}+\dot{y}_{r}.$
Note that (6) builds the relationship between state ${\boldsymbol{\xi}}$ and the control input ${\boldsymbol{u}}$.
@@
where $i=0,\cdots,H$, ${\boldsymbol{u}}_{k}=\{{\boldsymbol{u}}(k),\cdots,{\boldsymbol{u}}(k+H)\}$, $\omega_{r}^{\max}$, $\omega_{n}^{\max}$, and $v_{m}$ are the maximum velocity limits of the robot and nozzle driving motors and nozzle motion, respectively. We use YALMIP [43] to solve the MPC design in (V-C).

### v2

@@ 4 $\textit{conn}\leftarrow{\tt cell\_connection}({\boldsymbol{\Pi}}_{w},{{\boldsymbol{E}}})$
                  6 $\mathbb{G}_{c}\leftarrow{\tt Crack\_Graph}(\mathcal{I})$, ${\boldsymbol{\Pi}}_{c}\leftarrow{\tt GCC}(\mathbb{G}_{c})$
                  7 ${\tt follow\_path}({\boldsymbol{\Pi}}_{c}(i).\textit{nodes})$
                  8 $\mathcal{M}_{\text{cov}}\leftarrow\mathcal{M}_{\text{cov}}\cup(\mathbb{G}_{c}\oplus S)$
                   break
@@ 8 $\mathcal{M}_{\text{cov}}\leftarrow\mathcal{M}_{\text{cov}}\cup(\mathbb{G}_{c}\oplus S)$
            10 $\mathcal{M}_{\text{cov}}\leftarrow\mathcal{M}_{\text{cov}}\cup(P_{\text{n}}\oplus S),\;\mathcal{I}\leftarrow{\tt get\_image}$
            11 $\mathcal{T}\leftarrow{\tt get\_topology}(\mathcal{I})$
@@
In Algorithm 4, after traversing one target edge, we remove the covered areas, and update $\mathbb{G}_{\text{w}}$ of the remaining space $\mathcal{W}\backslash\mathcal{M}_{\text{cov}}$ to avoid passing the same target twice to reach another uncovered cell. Taking the example shown in Fig. 6, the robot is currently on node $A$ and $\mathbb{G}_{\text{w}}$ is updated after removing the already covered areas (yellow shaded areas). The odd nodes in the resulting $\mathbb{G}_{\text{w}}$ include $C_{1}$, $C_{2}$, $C_{3}$, and $C_{4}$. Therefore, $AC_{2}$ and $C_{1}C_{3}$ are connected to form the least-cost Euler tour, and $C_{4}$ represents the path’s ending node. When the resulting Euler tour needs to double edges in $\mathbb{G}_{\text{w}}$, the corresponding cell is split into two components. The first part is covered by a wall-following motion, where a wall is defined as the boundaries of $\mathcal{W}\backslash\mathcal{M}_{\text{cov}}$. The other part is covered by the zigzag motion of the leftover space in the cell. As shown in Fig. 6, edges $E_{1}$ and $E_{2}$ are doubled, and in Fig. 6, edges $E_{2}$ and $E_{3}$ are doubled. Note that the splitting of cells does not increase the cost of covering the whole cell. To minimize the cost, we select the coverage motion direction according to the next connected edge in the path.
@@
The ${\tt oSCC}$ algorithm ensures that each individual cell in the free space is covered exactly once, thereby avoiding redundancy in terms of individual cell coverage when the connections of each covered cell are not considered. However, due to the unknown dimensions of the cells, the robot may not always have enough information to minimize the zigzag motion connecting adjacent cells according to the Euler tour sequence. The optimality gap in our approach is primarily related to the connections between adjacent cells based on the Euler tour sequence. For worst-case analysis of the suboptimality bound, assuming that the connections between each pair of adjacent cells are small enough to be negligible in the optimal solution, the worst connection between each pair of adjacent cells resulting from the ${\tt oSCC}$ could be the number of free cells in $\mathbb{G}_{\text{w}}$, denoted as $N_{\text{cell}}$, minus one (i.e., the number of connections), multiplied by $\sqrt{l^{2}+(2S)^{2}}$ (i.e., the longest robot travel distance for one connection), where $l$ is the slice length (the side length of the free space). Therefore, in the worst case, the difference from the optimal solution is bounded by $(N_{\text{cell}}-1)\sqrt{l^{2}+(2S)^{2}}$. The simplified $\mathbb{G}_{\text{w}}$ reduces the number of edges. The algorithm is solved in polynomial time because of the structure of $\mathbb{G}_{\text{w}}$ and is therefore used efficiently online.
@@
Figure 7: Planning result by the ${\tt SCC}$ algorithm with the cracks as in Fig. 6. (a) The MCD and the corresponding $\mathbb{G}_{\text{w}}$ using ${\tt SCC}$. The orange dashed lines are the constructed $\mathbb{G}_{\text{c}}$. The blue-dotted dashed lines show the node connections. (b) The final route of the robot. The legends are the same as those in Fig. 6.
We further illustrate the ${\tt oSCC}$ and ${\tt SCC}$ planners through an example. Fig. 6 shows the planning result under ${\tt oSCC}$ in $\mathcal{W}$, with dimensions $l=5.79$ m and $w=6.10$ m, and robot configuration, $S=0.69$ m and $a=8.9$ cm. The crack information (thick, black, dotted lines) is initially unknown. Under the ${\tt oSCC}$ planner, the robot first follows the initial path planned by Algorithm 4 until detecting a crack node. Then it follows the online updated $\mathbb{G}_{\text{c}}$ and uses the ${\tt GCC}$ algorithm to generate local paths for the footprint coverage of the scanned cracks. The covered regions are then removed from $\mathcal{W}$, and $\mathbb{G}_{\text{c}}$ and $\mathbb{G}_{\text{w}}$ are updated accordingly. Figs. 6-6 show the updated graphs, and Fig. 6 illustrates the final path. For comparison, Fig. 7 presents the planning results under the ${\tt SCC}$ algorithm with known crack information. The constructed MCD of the free space is shown in Fig. 7, along with the corresponding $\mathbb{G}_{\text{w}}$ (solid lines), $\mathbb{G}_{\text{c}}$ (dash lines), and $\mathcal{M}_{\text{c}}$ (yellow shaded areas). The least-cost Euler tour is constructed by the blue dotted dash lines. Fig. 7 demonstrates the final route of the robot. For further details, readers can refer to the companion video clip. Experiments and comparisons will be discussed in Section VII.
@@ V-B Filling Nozzle Motion Planning
The nozzle motion needs to be coordinated with the robot’s motion to efficiently fill cracks within $\mathcal{F}$; see Fig. 8. Note that multiple cracks can be located within $\mathcal{F}$ at a time, and we need to determine how to move the nozzle to fill these cracks while the robot is in motion. Generally, the nozzle motion is much faster compared to the robot’s movement velocity. Therefore, for simplicity, we neglect the time duration for the nozzle to move from one crack to another without performing filling action and only consider the movement time along cracks during the filling action.
Fig. 9 illustrates the geometric relationship between the robot trajectory $\mathcal{P}_{R}$ and multiple cracks within $\mathcal{F}$. We use the arc length of the path $\mathcal{P}_{R}$, denoted as $s$, as the parameter to characterize any arbitrary point $p$ on the cracks. For a point $p$ on a crack within $\mathcal{F}$, we define a projection map $\pi(p):\;{p}\mapsto p_{c}(s)$, where $p_{c}(s)\in\mathcal{P}_{R}$, as the minimum distance to $p$. Considering that there are $n_{c}$ cracks located within $\mathcal{F}$, where $n_{c}\in\mathbb{N}$, we denote the mapping $\pi_{i}(p)$ for the $i$th crack by the above definition, where $i=1,\cdots,n_{c}$. We assume that $\pi_{i}(p)$ is bijective so that its inverse $\pi^{-1}_{i}(p)$ exists. By using $\pi_{i}(p)$, all points on the $n_{c}$ cracks inside $\mathcal{F}$ are mapped onto $\mathcal{P}_{R}$.
@@
Value $s$ increases along the robot $\mathbf{R}$’s moving direction, and points on $n_{c}$ cracks within $\mathcal{F}$ with the same $s$ value have the same priority for filling. To determine the filling sequences, we consider the nozzle $\mathbf{N}_{z}$ to move and switch the filling action among these $n_{c}$ cracks after staying along one crack for a threshold distance $L$. Fig. 9 illustrates the nozzle travel sequence among all cracks within $\mathcal{F}$. Algorithm 5 describes the nozzle motion sequence planning. Input $\{{\boldsymbol{p}}_{i}\}_{1}^{n_{c}}$ is the point sequence sets of all cracks; that is, $\{{\boldsymbol{p}}_{i}\}$ contains the point sequence of the $i$th crack, $i=1,\cdots,n_{c}$. ${\tt startpoint}$ (line 1) is to find the first point of each crack. The function ${\tt findmin}$ is to find the minimum value and the corresponding crack index. The notation $h(i)/h(I_{1})$ indicates removing $h(I_{1})$ from all $h(i)$ (line 3). The function ${\tt addpoint}$ (line 4) is to add an interval point of the crack $I_{1}$ to ${}^{s}{\boldsymbol{p}}_{m}$, and the interval is from $p_{1}$ to $p_{2}+L$. Finally, we obtain the output $\{{\boldsymbol{p}}_{n}\}$ that represents the point sequence of the nozzle in $\mathcal{B}$.
@@
1 ${\boldsymbol{p}}_{c}^{i}\leftarrow\pi_{i}({\boldsymbol{p}}_{i})$, $h(i)\leftarrow{\tt startpoint}({\boldsymbol{p}}_{c}^{i})$, $i=1,\cdots,n_{c}$
2 $(p_{1},I_{1})\leftarrow{\tt findmin}(h(i))$, $m\leftarrow 0$
while $m<n_{c}$ do
      3 $(p_{2},I_{2})\leftarrow{\tt findmin}[h(i)/h(I_{1})]$
      4 ${}^{s}{\boldsymbol{p}}_{m}\leftarrow{\tt addpoint}({\boldsymbol{p}}_{c}(I_{1}),p_{1},p_{2}+L)$
      5 if ${\tt isfinish}(I_{1})$ then $m\leftarrow m+1$, $h(I_{1})\leftarrow inf$
      6 else  $h(I_{1})\leftarrow p_{2}+L$
      7 $(p_{1},I_{1})\leftarrow(p_{2},I_{2})$
@@
As depicted in Fig. 9, according to Algorithm 5, the first points for each crack of the first three cracks are $h(1)=s_{1a}$, $h(2)=s_{2a}$, and $h(3)=s_{3a}$. In the algorithm, $p_{1}=s_{1a}$ and $I_{1}=1$ since $s_{1a}$ is the minimum point. Then, we find the minimum from $h(2)$ and $h(3)$, so that $p_{2}=s_{2a}$, $I_{2}=2$ (line 3). We also obtain $s_{1b}=s_{2a}+L$ (line 4), and therefore, all points on crack 1 from $s_{1a}$ to $s_{1b}$ are added to ${}^{s}{\boldsymbol{p}}_{1}$. The next iteration then adds the points on crack 2 and continues until all $n_{c}$ cracks are covered. The value of $L$ is chosen to provide a trade-off between the nozzle switching frequency among cracks and the dedicating time duration to a single crack. When $L$ is small, the switching between different cracks becomes frequent, and when $L$ is large, it might miss filling some cracks within $\mathcal{F}$. We consider using $L\leq 2a$ for the switching between cracks within $\mathcal{F}$ to have good performance.
@@ V-C Robot and Nozzle Motion Control
The robot’s orientation is independently controlled with linear velocity due to the use of omni-directional wheels. We first discuss the choice of orientation control. The robot’s driving energy expense can be represented as $J_{\theta}=\frac{1}{2}\sum_{i=1}^{4}\omega_{i}^{2}$. Taking the derivative of $J_{\theta}$ with respect to $\theta$, we obtain
@@
Using (1), it is straightforward to show that $\frac{\partial J_{\theta}}{\partial\theta}=0$, indicating that the energy expense is independent of the robot’s orientation. Similarly, we obtain
@@
The above result implies that changing the orientation increases the energy cost. Therefore, from (3) and (4), we set $\dot{\theta}=0$ to minimize the energy expense $J_{\theta}$, and in implementation, we further maintain $\theta=0$ as the desired body orientation during robot movement for simplicity.
For the robot and nozzle motion control, the objective is to reach all cracks within $\mathcal{F}$ and complete the crack-filling action while robot $\mathbf{R}$ is in motion. The desired nozzle path is given by Algorithm 5 in $\mathcal{B}$ as $\{{\boldsymbol{p}}_{n}\}$, while the desired robot path $\mathcal{P}_{R}$ is denoted as ${\boldsymbol{p}}_{r}$ in $\mathcal{N}$. $\mathcal{P}_{R}$ is computed from the coverage path planning algorithms, specifically Algorithms 3, 4, and 6. Defining ${\boldsymbol{\xi}}=[{\boldsymbol{q}}_{r}^{\mathsf{T}}\;{\boldsymbol{q}}_{n}^{\mathsf{T}}]^{\mathsf{T}}$ and ${\boldsymbol{u}}=[{\boldsymbol{\omega}}_{r}^{\mathsf{T}}\;{\boldsymbol{\omega}}_{n}^{\mathsf{T}}]^{\mathsf{T}}$, from (1), a discrete-time state-space model is used to represent the robot and nozzle motions at the $k$th step
@@
The robot velocity is slow compared to the nozzle motion. We denote $\mathcal{X}_{r}\subset\mathbb{R}^{3}$ and $\mathcal{X}_{n}\subset\mathbb{R}^{2}$ as the allowable robot and filling nozzle velocity sets in $\mathcal{N}$, respectively. We then have the velocity constraints as $\|{\boldsymbol{v}}_{r}(k)\|_{2}\leq\|{\boldsymbol{v}}_{n}(k)\|_{2}$, where ${\boldsymbol{v}}_{r}(k)=\dot{{\boldsymbol{q}}}_{r}(k)\in\mathcal{X}_{r}$ and ${\boldsymbol{v}}_{n}(k)=\dot{{\boldsymbol{q}}}_{n}(k)\in\mathcal{X}_{n},$ are the robot velocity and the nozzle relative velocity in $\mathcal{N}$, respectively. Using (1), ${\boldsymbol{v}}_{r}$ and ${\boldsymbol{v}}_{n}$ are calculated as
@@
$v_{n1}$ $=$ $\frac{R_{g}}{2}(\omega_{5}+\omega_{6})\mathop{\rm c}\nolimits_{\theta}-{x_{n}}\dot{\theta}\mathop{\rm s}\nolimits_{\theta}+\frac{R_{g}}{2}(\omega_{5}-\omega_{6})\mathop{\rm s}\nolimits_{\theta}-$
${y_{n}}\dot{\theta}\mathop{\rm c}\nolimits_{\theta}+\dot{x}_{r},$
$v_{n2}$ $=$ $\frac{R_{g}}{2}(\omega_{5}+\omega_{6})\mathop{\rm s}\nolimits_{\theta}+{x_{n}}\dot{\theta}\mathop{\rm c}\nolimits_{\theta}-\frac{R_{g}}{2}(\omega_{5}-\omega_{6})\mathop{\rm c}\nolimits_{\theta}-$
${y_{n}}\dot{\theta}\mathop{\rm s}\nolimits_{\theta}+\dot{y}_{r}.$
Note that (6) establishes the relationship between state ${\boldsymbol{\xi}}$ and the control input ${\boldsymbol{u}}$.
@@
where $i=0,\cdots,H$, ${\boldsymbol{u}}_{k}=\{{\boldsymbol{u}}(k),\cdots,{\boldsymbol{u}}(k+H)\}$, $\omega_{r}^{\max}$, $\omega_{n}^{\max}$, and $v_{m}$ are the maximum velocity limits of the robot and nozzle driving motors and nozzle motion, respectively. We use YALMIP [43] to solve the MPC design in (8).